OpenStudy (anonymous):

What is the GCF in this expression? 3y²(y - 2) + 5(y - 2)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

@Avon ???

OpenStudy (anonymous):

you can easily see which is G.C.F

OpenStudy (skullpatrol):

What happens if you multiply it out?

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (skullpatrol):

$$\huge \text{No effort = No answer}$$

OpenStudy (skullpatrol):

@Avon Well, are you ready to learn?

OpenStudy (anonymous):

yea @skullpatrol

OpenStudy (skullpatrol):

What happens if you multiply it out?

OpenStudy (anonymous):

@skullpatrol the site is acting crazy thts y i could answer, but 3y^3+5y^2 is tht right

OpenStudy (skullpatrol):

Yes, there is no common factor if you multiply it out.

OpenStudy (skullpatrol):

@surjithayer please don't give away the answer

OpenStudy (anonymous):

@skullpatrol so is 3 and 5 the GCF 3*1 AND 5*1 or is it 3y^2 =5 and that may equal 8y^2 do the y-2 cancel out

OpenStudy (skullpatrol):

$$\huge 3y²(y - 2) + 5(y - 2)$$

OpenStudy (anonymous):

3y+5y-2

OpenStudy (skullpatrol):

What is that?

OpenStudy (anonymous):

or is it 15=3*5

OpenStudy (skullpatrol):

Do you know what a "common factor" is"

OpenStudy (anonymous):

3+5=8

OpenStudy (skullpatrol):

What is a "factor"?

OpenStudy (anonymous):

3y²(y - 2) + 5(y - 2)= 3y^3-6y^2+5y-10= 1(3y^3-6y^2) + 1(5y-10)

OpenStudy (anonymous):

im not good at this idk seriously

OpenStudy (skullpatrol):

factor: When two or more numbers are multiplied each of the numbers is a factor of the product.

OpenStudy (anonymous):

ok i know that much @skullpatrol but what your not understanding is that im confused

OpenStudy (skullpatrol):

You need to recall the distributive property: $$\huge ba+ca=(b+c)a$$ Remember?

OpenStudy (skullpatrol):

$$\huge 3y²(y - 2) + 5(y - 2)=?$$

OpenStudy (skullpatrol):